OpenStudy (anonymous):

quadratic in vertex form y=x^2-14x+53

OpenStudy (anonymous):

do you know the steps for converting a quadratic to vertex form?

OpenStudy (anonymous):

half of \(-14\) is \(-7\) so write it as \[y=(x-7)^2+k\] so find \(k\) replace \(x\) by \(7\) in the original standard form, in other words \[7^2-14\times 7+53=k\]

OpenStudy (anonymous):

so when i am solving for x do i just forget about the 53

OpenStudy (anonymous):

you are not "solving for \(x\)" the question was to write is "vertex form" i.e. to turn \[y=x^2-14x+53\] in to \[y=(x-h)^2+k\]

OpenStudy (anonymous):

i know but i am talking using x to get the y so i can get the vertex

OpenStudy (anonymous):

\(h=\frac{b}{2a}\) which is this case is \(\frac{14}{2}=7\) so you start with \[y=(x-7)^2+k\]

OpenStudy (anonymous):

ok i get it

OpenStudy (anonymous):

then you need \(k\) which is \[7^2-14\times 7+53=49-98+53=4\]

OpenStudy (anonymous):

also i can show u another way to do it and its by factoring

OpenStudy (anonymous):

final answer is \[y=(x-7)^2+4\]

OpenStudy (anonymous):

could i see the way by factoring

OpenStudy (anonymous):